# Networking Systems for Video Anomaly Detection: A Tutorial and Survey

### v2

@@ 4.2. Classic WsVAD Models
We select two representative methods from unimodal and multimodal WsVAD, i.e., Multi-Instance Ranking (MIR) (Sultani et al., 2018) and Holistic and Localized Network (HL-Net) (Wu et al., 2020), to elaborate on the concepts of MIL and multimodal information processing in WsVAD research, respectively. Specifically, MIR has laid the foundation for the MIL-based solution, marking a milestone in VAD research. However, its optimization objectives and detailed implementation are challenging for many researchers in the AD and CV communities to understand. Therefore, we provide a comprehensive exposition of the motivation, theoretical logic, and related knowledge. In contrast, HL-Net is the first model for multimodal violence detection. The simultaneously released XD-violence dataset has inspired lots of researchers from the AD and multimodal understanding fields to delve into the emerging hotspot of multimodal VAD. We choose this method as a case study to disseminate knowledge of multimodal understanding to the AD community, aiming to propel multimodal VAD from simple-modal fusion towards systematic multimodal anomaly clues exploration, please see Section 2.2 of Supplement for more information.
@@ 6. Model Evaluation
Figure 7. Exapmles of classical UVAD (Subway and UMN) and WsVAD (UCF-Crime and XD-Violence) dataets.
In this section, we elaborate on the characteristics of prevailing datasets and common evaluation metrics. Based on the annotations, we categorize existing datasets into unsupervised, weakly supervised, and supervised, as presented in Table 3. The examples of UVAD (e.g., Subway and UMN) and WsVAD datasets are illustrated in Fig. 7. Since such datasets have been extensively surveyed, we only present their details in Section 4 of Supplement. We provide a systematic WSVAD evaluation system in this section, categorizing existing metrics into accuracy-oriented and cost-oriented as well as introducing system-level performance indicators.
@@ 6.1.1. Accuracy-oriented metrics
Due to the highly imbalanced nature of positive and negative samples in NSVAD, some common evaluation metrics for classification tasks, such as accuracy, are not applicable. For example, a model biased towards outputting label 0 thus missing anomalous events, would be incorrectly evaluated as good under such metrics. Using a single threshold to simply assess a model’s ability to differentiate between normal and abnormal patterns is not wise. For instance, with a threshold of 0.5, a model that consistently outputs scores slightly below 0.5 for regular events and slightly above 0.5 for anomalies would be considered perfect because it would show optimal performance across various metrics. However, such a model may not have learned the inherent differences between normal and abnormal patterns well, resulting in minimal score gaps between the two, which could lead to failure in detecting subtle anomalies and normal instances with data bias in complex scenarios. Therefore, researchers have introduced Receiver Operating Characteristic (ROC) curves, which measure VAD models more comprehensively by selecting multiple thresholds. Specifically, this curve plots the True Negative Rate (TNR) against the True Positive Rate (TPR) at various thresholds. The Area Under the ROC curve (AUROC) has been the most widely used VAD evaluation metric. An ideal AUROC value of 1 indicates a model that outputs a score of 0 for all negative samples and 1 for all positive samples, aligning with our expectations. Considering that TN usually exceeds TP, researchers argue that Average Precision, i.e., the Area Under the Precision-Recall (AUPR) curve, is more suitable for evaluating anomaly detection tasks. The PR curve depicts precision and recall (i.e., TPR) at specific thresholds. The point on this curve where Precision equals Recall is the balance point. Currently, multimodal anomaly detection models primarily use AP for quantitative evaluation. An anomaly score curve is commonly used to intuitively demonstrate the model’s response to anomalous events, presenting the temporal localization capability. In contrast, prediction error maps are often used to assess UVAD model’s spatial localization capability.
@@ 6.2. System-level Evaluation
While conventional metrics focus on the detection accuracy and cost of individual models, evaluating the performance of NSVAD in real-world deployments requires a more holistic, system-level perspective. Beyond model-specific metrics, the entire system’s effectiveness depends on various factors such as latency, communication cost, bandwidth efficiency, data security, user privacy, and system robustness. These metrics collectively assess the performance and feasibility of deploying NSVAD in large-scale, distributed, and resource-constrained environments. To this end, we categorize system-level metrics into three primary groups: efficiency, privacy, and robustness.
@@ 6.2.2. Privacy-oriented metrics
In NSVAD systems, maintaining data security and user privacy is paramount, particularly in applications involving sensitive environments such as healthcare (Li et al., 2024c) or public surveillance. Data security is typically evaluated using metrics such as encryption overhead, which measures the computational cost introduced by encryption techniques, and key management efficiency, which assesses how well the system handles the distribution and renewal of cryptographic keys in a large-scale deployment. Robust encryption algorithms, such as AES or homomorphic encryption, are commonly used to ensure that video data remains secure during transmission and processing. User privacy is another critical concern in video anomaly detection systems. Metrics such as the privacy leakage rate evaluate how much sensitive information (e.g., identities, personal activities) can be inferred from the system’s outputs. Differential privacy, federated learning, and encrypted video coding are among the techniques employed to minimize privacy risks. Privacy-preserving methods are evaluated based on the degree of anonymization they provide, often measured in terms of the privacy budget, which balances privacy protection against utility loss.
@@ 6.2.3. Robustness-oriented metrics
Robustness metrics aim to assess the system’s ability to maintain reliable performance in the presence of adversarial conditions, such as network disruptions, data corruption, or malicious attacks. One critical metric is fault tolerance, which measures the system’s capacity to continue operating when certain components fail. This is especially important in distributed settings where failures in edge devices or communication links can affect the overall detection pipeline. Techniques such as redundancy, dynamic task migration, and edge-cloud coordination can enhance the system’s fault tolerance. Adversarial robustness evaluates the system’s resilience to attacks designed to manipulate or mislead the anomaly detection model. Adversarial attacks may involve injecting malicious data, such as perturbing video frames or manipulating model parameters. The robustness against such attacks is typically quantified by the system’s ability to maintain high detection accuracy even when exposed to adversarial perturbations. Metrics such as adversarial success rate and robust accuracy (accuracy under attack) are often used to evaluate this aspect. Lastly, scalability measures the system’s ability to handle increasing workloads, including the number of video streams and distributed clients. This is typically evaluated through stress testing, where system performance is analyzed under different load conditions to ensure it can maintain efficiency and reliability as deployment scales up. A scalable NSVAD system should efficiently distribute workloads across edge devices and cloud servers without degrading performance.
7. Research Cases
To intuitively demonstrate the deployment potential and challenges faced by NSVAD in real-world applications, we selected three of our research cases to apply NSVAD in modern industry (Liu et al., 2024b), smart cities (Liu et al., 2023e), and large-scale systems.
7.1. Appearance-Motion Prototype Network for Smart Industry
NSVAD can capture global spatial-temporal information in open spaces and detect various anomalous events in industrial settings to ensure production safety and stability. Compared to traditional monitoring devices, video camera data provides stable quality and high information density, making NSVAD the preferred solution for Automatic Industrial Anomaly Detection Systems (AIADS). However, current VAD research lacks a deep focus on industrial scenarios, and existing methods often target outdoor public spaces, struggling with the high-temporal dynamics and complex spatial-temporal interactions in manufacturing environments. The scarcity of industrial video datasets limits AIADS development, while the diversity and complexity of industrial events make it challenging to develop effective NSVAD systems capable of representing regular patterns and generalizing to anomalies. Inspired by the UVAD anomaly categorization in Section 2.2 and the multi-stream approach in Section 3.1, we propose the first NSVAD model for AIADS, named AMP-Net, as shown in Fig. 9 (Liu et al., 2024b). Experiments on real-world industrial datasets show AMP-Net’s ability to handle dynamic environments while maintaining limited generalization to industrial anomalies.
Figure 9. Structure overview of the AMP-Net.
As the first NSVAD model for industrial scenarios, AMP-Net demonstrates the feasibility of using video for automated production monitoring and the potential of NSVAD in smart industries. Its development was influenced by earlier work on multi-stream approaches (Liu et al., 2022c) and memory networks (Gong et al., 2019; Park et al., 2020), which will inspire future researchers designing NSVAD systems for specific applications, as explained in Section 3.1 of Supplement. Although AMP-Net performs well on self-collected industrial datasets, it does not account for deployment costs on terminal devices or handle arbitrary data shifts in real-world environments, indicating further research is needed before it can be widely applied in modern factories. Additionally, analyzing the model’s detection capabilities for different industrial anomaly events and improving its detection of highly hazardous anomalies remain important goals.
7.2. Causal Video Normality Learning Towards Smart Cities
Figure 10. Pipeline of causal video normality learning.
Smart cities are reshaping lifestyles through AI and next-generation communication, relying on data sensing and understanding. Among the multimodal data from applications like modern transportation, smart homes, and public services, videos dominate in volume and information density, making automated video processing fundamental to smart city research. NSVAD, designed for smart cities, promises solutions for traffic accident detection (Liu et al., 2023c), elderly monitoring (Li et al., 2024b), and public safety. However, challenges like label-independent data shifts and subtle anomalies from event diversity remain. UVAD methods (Liu et al., 2018b; Gong et al., 2019; Chang et al., 2020), including AMP-Net (Liu et al., 2024b), have aimed to balance representation and generalization, but deep learning models still struggle to establish simple mappings between regular events and labels, lacking exploration of causal relationships in video regularity. Our case studies (Liu et al., 2023e) show that deep representation learning (DRL) often fails to distinguish normal and abnormal events, leading to performance degradation in complex scenarios.
To address this, we introduced Causal Representation Learning (CRL) into NSVAD, proposing Causal Video Normality Learning (CVNL) to explore causal relationships (Liu et al., 2023e). Unlike DRL’s statistical dependency-based description, CVNL identifies anomalies based on causal consistency in regular events, enabling better generalization to diverse regular events while remaining sensitive to anomalies. Extensive experiments on large-scale cross-scene datasets show that CVNL meets detection requirements in complex environments like smart cities, maintaining stable performance even with limited regular event data. The simplified CVNL framework is shown in Fig. 10, with more details in (Liu et al., 2023e) and Section 3.2 of Supplement.
7.3. End-to-Cloud Collaborative Framework for NSVAD
Inspired by NSAI (Song et al., 2022), we propose a hierarchical NSVAD framework for large-scale IoT and mobile internet applications, including hardware, system, algorithm, and application layers. Current NSVAD research focuses mainly on algorithms and applications, such as unsupervised crowd anomaly detection and multimodal violence detection. However, limited computational resources on terminal cameras restrict anomaly detection to short-term, localized events. For example, smart city traffic management requires large-scale systems to monitor traffic flows across a metropolitan area, but current methods only handle local detection at intersections or single devices.
While mobile communication allows devices like smartphones to use cloud servers for large-scale detection, raw data uploads significantly increase communication costs and raise privacy concerns. To address these challenges, we design a configurable end-to-cloud collaborative framework, balancing communication overhead and enhancing data security.
This framework enables terminal devices and cloud servers to collaborate on detecting various anomalies, including local, global, short-term, and long-term events. End devices, such as cameras or mobile devices, handle preprocessing and feature extraction, transmitting only feature data to the cloud, reducing sensitivity and privacy risks. The cloud aggregates features from multiple devices to build large models capable of global, long-term detection and offloads models or results back to the end devices as needed. The framework is flexible, allowing an unlimited number of devices to join and select execution tasks based on real-time performance and complexity, minimizing latency and communication costs while leveraging cloud computing for accurate detection without compromising data privacy.
Key components include computing containers, a deployment platform, and a data pipeline. The containers provide cross-platform, lightweight execution environments for NSVAD models with continuous updates. The deployment platform targets terminal devices like cameras and smartphones, locally deploying large models trained in the cloud. To optimize data transmission, we propose a lightweight data pipeline inspired by Alibaba’s Walle framework, reducing redundancy and communication overhead. The real-time channel uses optimized SSL and long-link communication to minimize latency. Case studies show that our framework reduces cloud workload by 78%, improves detection accuracy (frame-level AUC) by 12%, and lowers false positive rates by 46% compared to traditional cloud-based solutions.
@@ 8.1.1. Data
Real-world videos exhibit label-independent domain shifts due to variations in scenes, camera angles, and device configurations (Duong et al., 2023). These subtle differences in spatial-temporal patterns, while easily comprehended by humans, often lead to high false positive rates in NSVAD models (Liu et al., 2023e). Existing methods typically validate their models on datasets from a single scene to avoid this issue (Ramachandra et al., 2020). For instance, the UCSD dataset (Li et al., 2013) includes two distinct perspectives, but they are treated as separate datasets. The ShanghaiTech dataset (Liu et al., 2018b), despite spanning 13 scenes, is often treated as a single scene, leading to performance drops when compared to simpler datasets like UCSD Ped2 (Li et al., 2013) and CUHK Avenue (Lu et al., 2013). In real-world applications, such scene and device variations are inevitable, making it impractical to develop specialized models for every setup. To address this, some researchers (Acsintoae et al., 2022) have proposed using virtual engines to simulate anomaly events and generate richer positive samples. However, datasets like XD-Violence (Wu et al., 2020), which include movie and game scenes, differ from real-world anomalies, limiting the effectiveness of models trained on them. Bridging this domain gap between virtual and real anomalies is essential for deployable NSVAD systems. Additionally, multimodal NSVAD, while a growing field, remains confined to the fusion of RGB images and synchronized audio, neglecting novel modalities like language (Wu et al., 2024b) and texts (Pu et al., 2024), limiting its application in media streaming and broadcasting.
@@ 8.1.3. Model
Unsupervised NSVAD models benefit from self-supervised learning for spatial-temporal feature representation but face challenges with over-generalization, which can lead to missed anomalies (Park et al., 2020). The key challenge is balancing representation and generalization to reduce both false positives and negatives (Liu et al., 2024b). Approaches like memory networks and causal representation learning (Liu et al., 2023e) have shown promise, but performance on complex datasets remains inconsistent. In contrast, WsVAD can only detect predefined anomaly events, limiting its adaptability and failing to meet open-world requirements. While WsVAD models generally offer more reliable results than UVAD and FuVAD, their reliance on ranking loss is still debated. To advance NSVAD, there is a need for interpretable models and systems that integrate data encryption to address privacy concerns.
@@ 8.1.5. NSVAD-specialized
For real-world applications, NSVAD needs to handle vast data streams from urban surveillance cameras while addressing long-term anomaly detection. These systems also face public concerns regarding data overhead and privacy security. Unlike algorithmic research, NSVAD for large-scale applications presents three distinct challenges: (i) Efficient data exchange and task offloading (Dong et al., 2024) between millions of cameras and distributed servers. Current research focuses on enhancing algorithms but neglects the costs of data acquisition and transmission, which are critical in large-scale deployments. NSVAD must adopt new video compression (Gomes et al., 2023) and transmission protocols to handle the expanding network of cameras and increasing video resolution. Distributed machine learning techniques can also facilitate global model training without aggregating all data. (ii) Detecting anomalies across large spatial and temporal scales. Video IoT systems collect information from entire cities, but current NSVAD models (Su et al., 2025; Zhou et al., 2024), designed for discrete scenes, struggle with large-scale data. For example, traffic anomalies like congestion may be easily detected at intersections, but large-scale crowd movements might be misinterpreted as anomalous gatherings. (iii) Privacy protection and ethical concerns. NSVAD systems collect identifiable information such as faces and clothing, raising concerns about privacy breaches and biases related to race, gender, and skin color (Su et al., 2023, 2023; Noghre et al., 2024). While identity-agnostic data (e.g., encrypted video (Cheng et al., 2020)) and privacy-preserving techniques (e.g., federated learning (Al-Lahham et al., 2024)) are promising, ensuring transparency and fairness remains a significant challenge for NSVAD development.

### v3

@@ 4.2. Classic WsVAD Models
We select two representative methods from unimodal and multimodal WsVAD, i.e., MIR (Sultani et al., 2018) and HL-Net (Wu et al., 2020), to elaborate on the concepts of MIL and multimodal information processing in WsVAD research, respectively. Specifically, MIR has laid the foundation for the MIL-based solution, marking a milestone in VAD research. However, its optimization objectives and detailed implementation are challenging for many researchers in the AD and CV communities to understand. Therefore, we provide a comprehensive exposition of the motivation, theoretical logic, and related knowledge. In contrast, HL-Net is the first model for multimodal violence detection. The simultaneously released XD-violence dataset has inspired lots of researchers from the AD and multimodal understanding fields to delve into the emerging hotspot of multimodal VAD. We choose this method as a case study to disseminate knowledge of multimodal understanding to the AD community, aiming to propel multimodal VAD from simple-modal fusion towards systematic multimodal anomaly clues exploration, please see Section 2.2 of Supplement for more information.
@@ 6. Model Evaluation
Figure 7. Exapmles of classical UVAD (Subway (Adam et al., 2008) and UMN (Cui et al., 2011)) and WsVAD (UCF-Crime (Sultani et al., 2018) and XD-Violence (Wu et al., 2020)) dataets.
In this section, we elaborate on the characteristics of prevailing datasets and common evaluation metrics. Based on the annotations, we categorize existing datasets into unsupervised, weakly supervised, and supervised, as presented in Table 3. The examples of UVAD (e.g., Subway (Adam et al., 2008) and UMN (Cui et al., 2011)) and WsVAD (e.g., UCF-Crime (Sultani et al., 2018) and XD-Violence (Wu et al., 2020)) datasets are illustrated in Fig. 7. Since such datasets have been extensively surveyed, we only present their details in Section 4 of Supplement. We provide a systematic WSVAD evaluation system in this section, categorizing existing metrics into accuracy-oriented and cost-oriented as well as introducing system-level performance indicators.
@@ 6.1.1. Accuracy-oriented metrics
Due to the highly imbalanced nature of positive and negative samples in NSVAD, some common evaluation metrics for classification tasks, such as accuracy, are not applicable. For example, a model biased towards outputting label 0 thus missing anomalous events, would be incorrectly evaluated as good under such metrics. Using a single threshold to simply assess a model’s ability to differentiate between normal and abnormal patterns is not wise. For instance, with a threshold of 0.5, a model that consistently outputs scores slightly below 0.5 for regular events and slightly above 0.5 for anomalies would be considered perfect because it would show optimal performance across various metrics. However, such a model may not have learned the inherent differences between normal and abnormal patterns well, resulting in minimal score gaps between the two, which could lead to failure in detecting subtle anomalies and normal instances with data bias in complex scenarios. Therefore, researchers have introduced Receiver Operating Characteristic (ROC) curves, which measure VAD models more comprehensively by selecting multiple thresholds. Specifically, this curve plots the TNR against the TPR at various thresholds. The area under the curve, known as AUROC, has been the most widely used VAD evaluation metric. An ideal AUROC value of 1 indicates a model that outputs a score of 0 for all negative samples and 1 for all positive samples, aligning with our expectations. Considering that TN usually exceeds TP, researchers argue that Average Precision, i.e., the Area Under the Precision-Recall (AUPR) curve, is more suitable for evaluating anomaly detection tasks. The PR curve depicts precision and recall (i.e., TPR) at specific thresholds. The point on this curve where Precision equals Recall is the balance point. Currently, multimodal anomaly detection models primarily use AP for quantitative evaluation. An anomaly score curve is commonly used to intuitively demonstrate the model’s response to anomalous events, presenting the temporal localization capability. In contrast, prediction error maps are often used to assess UVAD model’s spatial localization capability.
@@ 6.2. System-level Evaluation
While conventional metrics focus on the detection accuracy and cost of individual models, evaluating the performance of NSVAD in real-world deployments requires a more holistic, system-level perspective. Beyond model-specific metrics, the entire system’s effectiveness depends on various factors such as latency, communication cost, bandwidth efficiency, data security, user privacy, and system robustness (Liu et al., 2024c). These metrics collectively assess the performance and feasibility of deploying NSVAD in large-scale, distributed, and resource-constrained environments. To this end, we categorize system-level metrics into three primary groups: efficiency, privacy, and robustness.
@@ 6.2.2. Privacy-oriented metrics
In NSVAD systems, maintaining data security and user privacy is paramount, particularly in applications involving sensitive environments such as healthcare (Li et al., 2024b) or public surveillance. Data security is typically evaluated using metrics such as encryption overhead, which measures the computational cost introduced by encryption techniques, and key management efficiency, which assesses how well the system handles the distribution and renewal of cryptographic keys in a large-scale deployment. Robust encryption algorithms, such as AES or homomorphic encryption, are commonly used to ensure that video data remains secure during transmission and processing. User privacy is another critical concern in video anomaly detection systems. Metrics such as the privacy leakage rate evaluate how much sensitive information (e.g., identities, personal activities) can be inferred from the system’s outputs. Differential privacy, federated learning, and encrypted video coding are among the techniques employed to minimize privacy risks. Privacy-preserving methods are evaluated based on the degree of anonymization they provide, often measured in terms of the privacy budget, which balances privacy protection against utility loss.
@@ 6.2.3. Robustness-oriented metrics
Robustness metrics aim to assess the system’s ability to maintain reliable performance in the presence of adversarial conditions, such as network disruptions, data corruption, or malicious attacks. One critical metric is fault tolerance, which measures the system’s capacity to continue operating when certain components fail (Guo et al., 2025b, a). This is especially important in distributed settings where failures in edge devices or communication links can affect the overall detection pipeline. Techniques such as redundancy, dynamic task migration, and edge-cloud coordination can enhance the system’s fault tolerance. Adversarial robustness evaluates the system’s resilience to attacks designed to manipulate or mislead the anomaly detection model. Adversarial attacks may involve injecting malicious data, such as perturbing video frames or manipulating model parameters. The robustness against such attacks is typically quantified by the system’s ability to maintain high detection accuracy even when exposed to adversarial perturbations. Metrics such as adversarial success rate and robust accuracy are often used. Lastly, scalability measures the system’s ability to handle increasing workloads, including the number of video streams and distributed clients. This is typically evaluated through stress testing, where system performance is analyzed under different load conditions to ensure it can maintain efficiency and reliability as deployment scales up. A scalable NSVAD system should efficiently distribute workloads across edge devices and cloud servers without degrading performance.
@@ 7.1.1. Data
Real-world videos exhibit label-independent domain shifts due to variations in scenes, camera angles, and device configurations (Duong et al., 2023). These subtle differences in spatial-temporal patterns, while easily comprehended by humans, often lead to high false positive rates in NSVAD models (Liu et al., 2023f). Existing methods typically validate their models on datasets from a single scene to avoid this issue (Ramachandra et al., 2020). For instance, the UCSD dataset (Li et al., 2013) includes two distinct perspectives, but they are treated as separate datasets. The ShanghaiTech dataset (Liu et al., 2018b), despite spanning 13 scenes, is often treated as a single scene, leading to performance drops when compared to simpler datasets like UCSD Ped2 (Li et al., 2013) and CUHK Avenue (Lu et al., 2013). In real-world applications, such scene and device variations are inevitable, making it impractical to develop specialized models for every setup. To address this, some researchers (Acsintoae et al., 2022) have proposed using virtual engines to simulate anomaly events and generate richer positive samples. However, datasets like XD-Violence (Wu et al., 2020), which include movie and game scenes, differ from real-world anomalies, limiting the effectiveness of models trained on them. Bridging this domain gap between virtual and real anomalies is essential for deployable NSVAD systems. Additionally, multimodal NSVAD, while a growing field, remains confined to the fusion of RGB images and synchronized audio, neglecting novel modalities like language (Wu et al., 2024b) and texts (Pu et al., 2024), limiting its application in media streaming and broadcasting.
@@ 7.1.3. Model
Unsupervised NSVAD models benefit from self-supervised learning for spatial-temporal feature representation but face challenges with over-generalization, which can lead to missed anomalies (Park et al., 2020). The key challenge is balancing representation and generalization to reduce both false positives and negatives (Liu et al., 2024d). Approaches like memory networks and causal representation learning (Liu et al., 2023f) have shown promise, but performance on complex datasets remains inconsistent. In contrast, WsVAD can only detect predefined anomaly events, limiting its adaptability and failing to meet open-world requirements. While WsVAD models generally offer more reliable results than UVAD and FuVAD, their reliance on ranking loss is still debated. To advance NSVAD, there is a need for interpretable models and systems that integrate data encryption to address privacy concerns.
@@ 7.1.5. NSVAD-specialized
For real-world applications, NSVAD needs to handle vast data streams from urban surveillance cameras while addressing long-term anomaly detection. These systems also face public concerns regarding data overhead and privacy security. Unlike algorithmic research, NSVAD for large-scale applications presents three distinct challenges: (i) Efficient data exchange and task offloading (Dong et al., 2024) between millions of cameras and distributed servers. Current research focuses on enhancing algorithms but neglects the costs of data acquisition and transmission, which are critical in large-scale deployments. NSVAD must adopt new video compression (Gomes et al., 2023) and transmission protocols to handle the expanding network of cameras and increasing video resolution. Distributed machine learning techniques can also facilitate global model training without aggregating all data. (ii) Detecting anomalies across large spatial and temporal scales. Video IoT systems collect information from entire cities, but current NSVAD models (Su et al., 2025; Zhou et al., 2024a), designed for discrete scenes, struggle with large-scale data. For example, traffic anomalies like congestion may be easily detected at intersections, but large-scale crowd movements might be misinterpreted as anomalous gatherings. (iii) Privacy protection and ethical concerns. NSVAD systems collect identifiable information such as faces and clothing, raising concerns about privacy breaches and biases related to race, gender, and skin color (Su et al., 2023, 2023; Noghre et al., 2024). While identity-agnostic data (e.g., encrypted video (Cheng et al., 2020)) and privacy-preserving techniques (e.g., federated learning (Al-Lahham et al., 2024)) are promising, ensuring transparency and fairness remains a significant challenge for NSVAD development.